Having a disagreement online. Does 6÷2(1+2) = 9 or 1? Thanks
Answer by stanbon(75887)

6÷2(1+2) = 9 or 1? Thanks
-----
The neumonic for the Rule: PEMDAS
P = Evaluate parentheses
E = Apply exponents
M/D : Multiply/Divide from left to right
A/S : Add/Subtract from left to right
=======================
= 6/2*3
----
= 3*3
===
= 9
===========
Cheers,
Stan H.